What is the angular displacement of a 0.7m diameter wheel that rolls 3m clockwise?

1
Expert's answer
2022-01-11T11:01:03-0500

The arc traveled is given by

"S=\theta\;r"

Where.

"S=3\;m" is the linear offset or arc traveled.

"r" Is the radius.

"\theta" Is the angular displacement.


The diameter is "d=0.7\;m"


The radius is half the diameter that is to say.

"r=\dfrac{d}{2}\\
r=\dfrac{0.7\;m}{2}\\
r=0.35\;m"


Obtaining the expression for the angular displacement.

"S=\theta\;r\\
\theta\;r=S\\
\theta=\dfrac{S}{r}\\"

Evaluating numerically.

"\theta=\dfrac{S}{r}\\
\theta=\dfrac{3.0\;m}{0.35\;m}\\
\theta=8.6\;rad"


The angular displacement is "\theta=8.6\;rad"